★ Fill in the blanks with correct word(s).



1 The volume of the gas would become zero at _______ temperature.



2 At constant temperature when the volume of the gas decreases, the pressure ________.



3. A sample of helium gas occupies 6 mL at a temperature of 250 K. At _______K the gas expands to 9 mL.



4. The pressure of a gas mixture is equal to the ________ of the partial pressures of the constituent gases at a particular temperature.



5. The value of PV for 5.6 moles of a gas at 0°C is ________.

1
Expert's answer
2022-01-31T20:05:01-0500

1)


Volume of gas become zero at "0K"


2)

At constant temperature, when volume of gas decreases, pressure increases.


3)

If pressure is constant

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{6}{250}=\frac{9}{T_2}"


"T_2=375K"


4)

Pressure of a gas mixture is equal to sum of the partial pressure of constituent gases.


5)

"PV=nRT"

"=5.6\u00d78.314\u00d7273"

=12710J